The first three terms of an arithmetic progression are 5x, 2x+3, and 2x+7 respectively. Find the value of x


1
Expert's answer
2021-02-10T14:29:43-0500

"a_1=5x"

"a_1+d=2x+3"

"a_1+2d=2x+7"


Substituting for the value "a_1"


"5x+d=2x+3"

"5x+2d=2x+7"


Forming simultaneous equation and solving by substitution ;


"3x+d=3"

"3x+2d=7"


"d=3-3x"

"6-6x+3x=7"

"6-3x=7"

"3x=13"


"x=\frac{13}{3}"